OpenStudy (anonymous):

Candy and Tim share a paper route. It takes Candy 95 min to deliver all the papers, and it takes Tim 55 min. How long does it take the two when they work together?

OpenStudy (anonymous):

150? idk

OpenStudy (anonymous):

Given that Cindy takes 95 min to complete the route and Tim takes 55 min to complete the route. Meaning in 1 min, Cindy complete 1/95 of the job while Tim completes 1/55 of the job. In 1 min, they complete 1/95 + 1/55 = 6/209 of the route. To find out how long they take to complete the whole route simply, take 209/6 = 34.8 mins approx.